OpenStudy (anonymous):

Convert. 45 mi/h = _____ ft/s

OpenStudy (anonymous):

@emily_wilson

OpenStudy (anonymous):

45 mph = 66 fps (feet per second)

OpenStudy (anonymous):



OpenStudy (anonymous):

do you get that?

OpenStudy (anonymous):

divide not side in the picture*

OpenStudy (anonymous):

yea so the answer is 66?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

OKAY

OpenStudy (anonymous):

THANK YOU SO MUCH AGAIN @emily_wilson

OpenStudy (anonymous):

srry i was on caps an didnt know

OpenStudy (anonymous):

You're Welcome haha!!

OpenStudy (anonymous):

=)

OpenStudy (mathstudent55):

Emily is correct. Write each conversion factor as a fraction, making sure you cancel out the units you don't want leaving the units you want. \(\large 45 \dfrac{\cancel{mi}}{\cancel{h}} \times \dfrac{5280~ft}{\cancel{mi}} \times \dfrac{\cancel{h}}{60~\cancel{mn}} \times \dfrac{\cancel{mn}}{60~s} = 66\dfrac{ft}{s}\)

OpenStudy (anonymous):

okay thnx